# Differentially Private Covariate Balancing Causal Inference

### v1

@@ 3.5 Variance Estimator
[…]
@@
Suppose we use the CBSR defined by (3) and (4) with $-1\leq\alpha,\beta\leq 0$. The $\ell_{1}$-sensitivity of $\hat{V}_{\alpha,\beta}$ and $\tilde{V}^{(1)}_{\alpha,\beta}$ is $\Delta_{V,\alpha,\beta}=1/2n\eta C_{\eta,\alpha,\beta}$ where $C_{\eta,\alpha,\beta}=\min\{\eta^{\alpha+1}(1-\eta)^{\beta+1},(1-\eta)^{\alpha+1}\eta^{\beta+1}\}$.
@@
Input: Database $D=\{(X_{i},Z_{i},Y_{i})\}_{i=1}^{n}$, positivity bound $\eta$, privacy budget $\epsilon$, privatized propensity score $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ from Algorithm 1
Output: Differentially private variance estimator $\tilde{V}^{*}_{\alpha,\beta}$
  1. 1.
    Compute the non-private estimator (14) by plugging in the propensity score $\hat{e}(x)$ obtained by solving (2).
  2. 2.
    Compute the partially privatized estimator $\tilde{V}^{(1)}_{\alpha,\beta}$ by plugging in $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(x)$ for $\hat{e}(x)$ in (14).
  3. 3.
    Privatize the variance estimator by applying the Laplace mechanism: $\tilde{V}^{(2)}_{\alpha,\beta}=\tilde{V}^{(1)}_{\alpha,\beta}+\nu_{V}$ where $\nu_{V}\sim\text{Lap}\left(0,\frac{\Delta_{V,\alpha,\beta}}{\epsilon}\right)$ and $\Delta_{V,\alpha,\beta}$ is chosen from Table 2 for the estimand of interest.
  4. 4.
    Output the following estimator:
    $\tilde{V}^{*}_{\alpha,\beta}=\begin{cases}\tilde{V}^{(2)}_{\alpha,\beta}&\text{if }\tilde{V}^{(2)}_{\alpha,\beta}>0\\ \frac{1}{4n\eta C_{\eta,\alpha,\beta}}+\frac{1}{2\epsilon^{2}n^{2}\eta^{2}}&\text{if }\tilde{V}^{(2)}_{\alpha,\beta}\leq 0\end{cases}$
@@
Since both $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$ separately satisfy $\epsilon$-DP, to obtain $\epsilon$-DP for the joint release, we need to scale the privacy budgets accordingly, but we do not need to allocate an equal budget to both $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$. Let $10<r<$ 1 denote the budget proportion for $\tilde{V}^{*}_{\alpha,\beta}$. We then apply the privacy mechanisms for $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$ such that they satisfy $(1-r)\epsilon$-DP and $r\epsilon$-DP respectively. Notably, we can reuse $\tilde{\theta}^{(1)}_{\alpha,\beta}$ for privatizing $\hat{V}^{*}_{\alpha,\beta}$, therefore, it is recommended that we allocate a smaller budget on $\tilde{V}^{*}_{\alpha,\beta}$, i.e., $r<0.5$. We discuss the choice of these hyperparameters in the supplementary material Section D.
@@ 4.3 Scenario 2: Mispecified Propensity Score Model
Table 3: Evaluation metrics under the well-specified scenario for our differentially private methodology versus the method of Guha and Reiter (2024) and Lee et al. (2019) under various sample sizes $N\in\{5000,10000,50000,100000\}$ and privacy budgets $\epsilon\in\{0.5,1.0,5.0\}$. The Lee et al. (2019)’s method does not offer the interval estimator, thus the coverage and interval length are not displayed.
MSE Bias Coverage Interval Length $N$ $\epsilon$ Our Method GR LGPM Our Method GR LGPM Our Method GR Our Method GR 0.5 $0.01425$ $0.02646$ $0.01266$ $0.41543$ $0.53978$ $0.39716$ $95.3\%$ $94.7\%$ $0.70084$ $0.65664$ 5000 1.0 $0.00378$ $0.00692$ $0.00625$ $0.21576$ $0.27788$ $0.29859$ $97.0\%$ $98.0\%$ $0.49685$ $0.46435$ 5.0 $0.00043$ $0.00067$ $0.00477$ $0.07581$ $0.09274$ $0.28739$ $97.0\%$ $100.0\%$ $0.22206$ $0.20795$ 0.5 $0.00394$ $0.01557$ $0.00556$ $0.21269$ $0.39357$ $0.29409$ $95.3\%$ $95.7\%$ $0.48965$ $0.54914$ 10000 1.0 $0.00110$ $0.00412$ $0.00447$ $0.11445$ $0.20580$ $0.27928$ $98.7\%$ $98.0\%$ $0.34867$ $0.38823$ 5.0 $0.00019$ $0.00036$ $0.00423$ $0.05023$ $0.06580$ $0.28273$ $97.7\%$ $100.0\%$ $0.15619$ $0.17357$ 0.5 $0.00016$ $0.00229$ $0.00398$ $0.04670$ $0.16421$ $0.28048$ $99.0\%$ $100.0\%$ $0.21604$ $0.36941$ 50000 1.0 $0.00007$ $0.00062$ $0.00391$ $0.02942$ $0.08585$ $0.28036$ $99.3\%$ $100.0\%$ $0.15383$ $0.26131$ 5.0 $0.00003$ $0.00007$ $0.00399$ $0.02117$ $0.03048$ $0.28413$ $96.0\%$ $100.0\%$ $0.06996$ $0.11668$ 0.5 $0.00005$ $0.00136$ $0.00383$ $0.02373$ $0.11743$ $0.27863$ $99.0\%$ $99.7\%$ $0.15522$ $0.30914$ 100000 1.0 $0.00002$ $0.00036$ $0.00384$ $0.01679$ $0.06168$ $0.27962$ $99.7\%$ $100.0\%$ $0.11052$ $0.21858$ 5.0 $0.00001$ $0.00004$ $0.00385$ $0.01380$ $0.02188$ $0.28008$ $99.0\%$ $100.0\%$ $0.04785$ $0.09769$
MSE, mean squared error; GR, the method of Guha and Reiter (2024); LGPM, the method of Lee et al. (2019).
@@
MSE Bias Coverage Interval Length $N$ $\epsilon$ Our Method GR LGPM Our Method GR LGPM Our Method GR Our Method GR 0.5 $0.01682$ $0.02637$ $0.06417$ $0.33168$ $0.38982$ $0.78137$ $93.7\%$ $22.0\%$ $0.69825$ $0.05990$ 5000 1.0 $0.00421$ $0.00681$ $0.05776$ $0.16577$ $0.20005$ $0.77720$ $97.0\%$ $98.0\%$ $0.49834$ $0.46435$ 5.0 $0.00030$ $0.00041$ $0.05686$ $0.04498$ $0.05331$ $0.78577$ $99.0\%$ $100.0\%$ $0.23289$ $0.20762$ 0.5 $0.00440$ $0.01555$ $0.05687$ $0.16513$ $0.28418$ $0.77599$ $94.7\%$ $19.3\%$ $0.49029$ $0.03619$ 10000 1.0 $0.00112$ $0.00391$ $0.05606$ $0.08423$ $0.14648$ $0.78001$ $97.7\%$ $98.7\%$ $0.34817$ $0.38824$ 5.0 $0.00011$ $0.00023$ $0.05629$ $0.02701$ $0.03898$ $0.78499$ $98.0\%$ $100.0\%$ $0.16101$ $0.17373$ 0.5 $0.00017$ $0.00219$ $0.05596$ $0.03481$ $0.11630$ $0.78359$ $99.3\%$ $13.7\%$ $0.21625$ $0.01217$ 50000 1.0 $0.00006$ $0.00058$ $0.05587$ $0.01983$ $0.06094$ $0.78346$ $99.7\%$ $100.0\%$ $0.15357$ $0.26129$ 5.0 $0.00002$ $0.00004$ $0.05583$ $0.01062$ $0.01701$ $0.78343$ $99.0\%$ $100.0\%$ $0.07037$ $0.11691$ 0.5 $0.00005$ $0.00134$ $0.05546$ $0.01717$ $0.08593$ $0.78102$ $99.7\%$ $12.3\%$ $0.15604$ $0.00775$ 100000 1.0 $0.00002$ $0.00035$ $0.05555$ $0.01058$ $0.04385$ $0.78172$ $99.7\%$ $100.0\%$ $0.11035$ $0.21863$ 5.0 $0.00001$ $0.00003$ $0.05563$ $0.00738$ $0.01310$ $0.78230$ $99.7\%$ $100.0\%$ $0.05180$ $0.09784$
MSE, mean squared error; GR, the method of Guha and Reiter (2024); LGPM, the method of Lee et al. (2019).
@@ 5 Data Analysis
For the PSID data, the point estimators are reasonably accurate, but because the PSID dataset is observational, the results are not as robust as those for the NSW dataset. The non-private CBSR estimate for the PSID data is $0.164$, with a confidence interval from $-0.075$ to $0.403$. Under privacy constraints, the estimates, such as the private CBSR with $\epsilon=1.0$, showing $0.035$ with a confidence interval of $-0.112$ to $0.182$, reflect greater variability. This variability emphasizes the trade-off between privacy and precision, particularly in observational settings where the underlying data characteristics are less controlled than in experimental settings like the NSW data.
@@
Results for NSW data Results for PSID data Estimator Estimate Confidence Interval Estimate Confidence Interval Non-private CBSR 0.112 $(0.023,0.201)$ 0.164 $(-0.075,0.403)$ Private CBSR ($\epsilon=0.5$) 0.067 $(-0.004,0.138)$0.007 $(-0.147,0.132)$ Private CBSR ($\epsilon=1.0$) 0.082 $(0.003,0.162)$ 0.035 $(-0.112,0.182)$ Private CBSR ($\epsilon=5.0$) 0.103 $(0.017,0.188)$ 0.122 $(-0.030,0.274)$
@@ 6 Concluding Remarks
A promising avenue for future research is the development of an analytical framework for unbounded variables. Our current framework is restricted to bounded variables due to the sensitivity considerations of differential privacy mechanisms. Additionally, the finite-sample performance of our estimators could potentially be improved by more carefully designing the second stage of our privatization process, allowing for a more tailored noise distribution (Awan and Slavković, 2021).
@@ A.1 Proof of Lemma 1
As discussed in Section 2, we consider a scenario where each unit $i$ has an outcome $Y_{i}\in[0,1]$, treatment assignment $Z_{i}\in\{0,1\}$, and covariates $X_{i}\in\mathcal{X}$, where $\mathcal{X}$ is the unit ball so that $\|X_{i}\|_{2}\leq 1$, respectively. Along with the boundedness assumptions, Proposition 2 of Barber and Duchi [2014] implies the minimax risk is lower bounded $\inf_{M_{\epsilon}\in\mathcal{M}_{\epsilon}}\inf_{\hat{\tau}}\sup_{P\in\mathcal{P}}\mathrm{E}\left\{(\hat{\tau}-\tau)^{2}\right\}=\Omega\left\{n^{-1}+(\epsilon n)^{-2}\right\}$. ∎
@@ A.2 Proof of Lemma 2
By Zhao [2019, Theorem 1], for $j=1,\ldots,d$, we have
@@ A.4 Proof of Proposition 3
We first introduce the technical assumptions in Hirano et al. [2003].
@@
The support of $X$ is a Cartesian product of compact intervals. The density of $X$ is bounded, and bounded away from 00.
@@
The convergence of the first component $\xi_{1}$ can be proved by applying the proof of Hirano et al. [2003], Zhao [2019]. Under Assumptions 16, Hirano et al. [2003], Zhao [2019] showed that $\hat{\tau}$ is a semiparametric efficient estimator for $\tau$ with the convergence rate $O_{p}(1/\sqrt{n})$. Thus $\xi_{1}$ converges to zero in probability and $\xi_{1}=O_{p}(1/\sqrt{n})$.
Next, let us consider $\xi_{2}$. We write $\tau(w)=\frac{\sum_{i=1}^{n}Z_{i}w(X_{i},1)Y_{i}}{\sum_{i=1}^{n}Z_{i}w(X_{i},1)}$, $\hat{w}=w(\hat{\theta})$ and $\tilde{w}=w(\tilde{\theta})$ to clarify the dependence. By the mean value theorem, there exists some $w_{0}$ such that
$\tilde{\tau}^{(1)}-\hat{\tau}=\tau(\tilde{w})-\tau(\hat{w})=\tau^{\prime}(w_{0})(\tilde{w}-\hat{w}),$
@@
$\tilde{w}-\hat{w}=w(\tilde{\theta})-w(\hat{\theta})=w^{\prime}(\theta_{0})(\tilde{\theta}-\hat{\theta}).$
Under Assumption 1, $\tau^{\prime}(w_{0})$ and $w^{\prime}(\theta_{0})$ are bounded. Additionally, according to Reimherr and Awan [2019, Theorem 3.2], the KNG mechanism ensures that, if the loss function is a twice differentiable convex loss function and the Hessian has strictly positive eigenvalues, we have $\tilde{\theta}-\hat{\theta}=O_{p}(1/np\epsilon).$ Zhao [2019, Proposition 2] showed this condition holds for $\alpha=-1,\beta=0$ or $\alpha=0,\beta=-1$ in the CBSR framework. Therefore, putting these together, we have
@@
=$=\{\tau+O_{p}(1/np\epsilon+1/\sqrt{n})+\nu_{1}/a_{n}\}\{1-\nu_{2}/a_{n}+(\nu_{2}/a_{n})^{2}-\cdots\}$

### v2

@@ 3.5 Variance Estimator
where $v_{z}(x)=\mathrm{var}\{Y(z)\mid X=x\}$ and $H_{\alpha,\beta}=\int h_{\alpha,\beta}(x)f(x)d\mu(x)$ is a normalizing constant. (13) can be estimated by the following estimator:
@@
Suppose we use the CBSR defined by (3) and (4) with $1\leq\alpha,\beta\leq 0$. The $\ell_{1}$-sensitivity of $\hat{V}_{\alpha,\beta}$ and $\tilde{V}^{(1)}_{\alpha,\beta}$ is $\Delta_{V,\alpha,\beta}=1/2n\eta C_{\eta,\alpha,\beta}$ where $C_{\eta,\alpha,\beta}=\min\{\eta^{\alpha+1}(1-\eta)^{\beta+1},(1-\eta)^{\alpha+1}\eta^{\beta+1}\}$.
@@
0: Database $D=\{(X_{i},Z_{i},Y_{i})\}_{i=1}^{n}$, positivity bound $\eta$, privacy budget $\epsilon$, privatized propensity score $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ from Algorithm 1.
0: Differentially private variance estimator $\tilde{V}^{*}_{\alpha,\beta}$.
1: Compute the non-private estimator by plugging in the propensity score $\hat{e}(x)$ obtained by solving the optimization equation.
2: Compute the partially privatized estimator $\tilde{V}^{(1)}_{\alpha,\beta}$ by plugging in $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(x)$ for $\hat{e}(x)$ in the non-private variance estimator.
3: Apply the Laplace mechanism to privatize the variance estimator:
$\tilde{V}^{(2)}_{\alpha,\beta}=\tilde{V}^{(1)}_{\alpha,\beta}+\nu_{V}$
where $\nu_{V}\sim\text{Lap}\left(0,\frac{\Delta_{V,\alpha,\beta}}{\epsilon}\right)$, and $\Delta_{V,\alpha,\beta}$ is selected based on the estimand of interest from Table 2.
4: Apply post-processing to ensure positivity of the variance estimator:
$\tilde{V}^{*}_{\alpha,\beta}=\begin{cases}\tilde{V}^{(2)}_{\alpha,\beta}&\text{if }\tilde{V}^{(2)}_{\alpha,\beta}>0,\\ \frac{1}{4n\eta C_{\eta,\alpha,\beta}}+\frac{1}{2\epsilon^{2}n^{2}\eta^{2}}&\text{if }\tilde{V}^{(2)}_{\alpha,\beta}\leq 0.\end{cases}$
5:return  $\tilde{V}^{*}_{\alpha,\beta}$.
@@
Remark (Choice of hyperparameters).
Since both $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$ separately satisfy $\epsilon$-DP, to obtain $\epsilon$-DP for the joint release, we need to scale the privacy budgets accordingly, but we do not need to allocate an equal budget to both $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$. Let $0<r<1$ denote the budget proportion for $\tilde{V}^{*}_{\alpha,\beta}$. We then apply the privacy mechanisms for $\tilde{\tau}_{\alpha,\beta}^{(2)}$ and $\tilde{V}^{*}_{\alpha,\beta}$ such that they satisfy $(1-r)\epsilon$-DP and $r\epsilon$-DP respectively. Notably, we can reuse $\tilde{\theta}^{(1)}_{\alpha,\beta}$ for privatizing $\hat{V}^{*}_{\alpha,\beta}$, therefore, it is recommended that we allocate a smaller budget on $\tilde{V}^{*}_{\alpha,\beta}$, i.e., $r<0.5$.
Specifically, our methodology involves three key hyperparameters: $p\in(0,1)$, $0<q_{j}<1$ for $j=1,\ldots,4$ with $\sum_{j=1}^{4}q_{j}=1$, and $r\in(0,1)$. The parameters $p$ and $q_{j}$ govern the privacy level of the point estimator, specifically controlling the allocation of the privacy budget between the first and second stages of privatization. When there is no prior information to guide the choice of these parameters, it is generally recommended to assign equal weights to each privatization. Specifically, we set $p=0.5$ and $q_{1}=\ldots=q_{4}=0.25$ as the default choice.
Finally, the positivity bound $\eta$ affects the estimators. It enters the variance estimator and the confidence intervals directly, whereas the point estimator is less sensitive. In our illustrative analysis, we use $\eta=0.05$ as the default. For applications in practice, we recommend assessing sensitivity to alternative choices of $\eta$.
@@ 4.3 Scenario 2: Mispecified Propensity Score Model
Table 3: Evaluation metrics under the well-specified scenario for our differentially private methodology versus the method of Guha and Reiter (2024) and Lee et al. (2019) under various sample sizes $N\in\{5000,10000,50000,100000\}$ and privacy budgets $\epsilon\in\{0.5,1.0,5.0\}$. The Lee et al. (2019)’s method does not offer the interval estimator, thus the coverage and interval length are not displayed. GR and LGPM represent the method of Guha and Reiter (2024) and Lee et al. (2019), repectively.
MSE Bias Coverage Interval Length NN $\epsilon$ Our Method GR LGPM Our Method GR LGPM Our Method GR Our Method GR 0.5 $0.01425$ $0.02646$ $0.01266$ $0.41543$ $0.53978$ $0.39716$ $95.3\%$ $94.7\%$ $0.70084$ $0.65664$ 5000 1.0 $0.00378$ $0.00692$ $0.00625$ $0.21576$ $0.27788$ $0.29859$ $97.0\%$ $98.0\%$ $0.49685$ $0.46435$ 5.0 $0.00043$ $0.00067$ $0.00477$ $0.07581$ $0.09274$ $0.28739$ $97.0\%$ $100.0\%$ $0.22206$ $0.20795$ 0.5 $0.00394$ $0.01557$ $0.00556$ $0.21269$ $0.39357$ $0.29409$ $95.3\%$ $95.7\%$ $0.48965$ $0.54914$ 10000 1.0 $0.00110$ $0.00412$ $0.00447$ $0.11445$ $0.20580$ $0.27928$ $98.7\%$ $98.0\%$ $0.34867$ $0.38823$ 5.0 $0.00019$ $0.00036$ $0.00423$ $0.05023$ $0.06580$ $0.28273$ $97.7\%$ $100.0\%$ $0.15619$ $0.17357$ 0.5 $0.00016$ $0.00229$ $0.00398$ $0.04670$ $0.16421$ $0.28048$ $99.0\%$ $100.0\%$ $0.21604$ $0.36941$ 50000 1.0 $0.00007$ $0.00062$ $0.00391$ $0.02942$ $0.08585$ $0.28036$ $99.3\%$ $100.0\%$ $0.15383$ $0.26131$ 5.0 $0.00003$ $0.00007$ $0.00399$ $0.02117$ $0.03048$ $0.28413$ $96.0\%$ $100.0\%$ $0.06996$ $0.11668$ 0.5 $0.00005$ $0.00136$ $0.00383$ $0.02373$ $0.11743$ $0.27863$ $9.0\% 99$$9.7\%$ $0.15522$ $0.30914$ 100000 1.0 $0.00002$ $0.00036$ $0.00384$ $0.01679$ $0.06168$ $0.27962$ $99.7\%$ $0.0\% 0.110$520.11052 $0.21858$ 5.0 $0.00001$ $0.00004$ $0.00385$ $0.01380$ $0.02188$ $0.28008$ $99.0\%$ $100.0\%$ $0.04785$ $0.09769$
@@
MSE Bias Coverage Interval Length NN $\epsilon$ Our Method GR LGPM Our Method GR LGPM Our Method GR Our Method GR 0.5 $0.01682$ $0.02637$ $0.06417$ $0.33168$ $0.38982$ $0.78137$ $93.7\%$ $22.0\%$ $0.69825$ $0.05990$ 5000 1.0 $0.00421$ $0.00681$ $0.05776$ $0.16577$ $0.20005$ $0.77720$ $97.0\%$ $98.0\%$ $0.49834$ $0.46435$ 5.0 $0.00030$ $0.00041$ $0.05686$ $0.04498$ $0.05331$ $0.78577$ $99.0\%$ $100.0\%$ $0.23289$ $0.20762$ 0.5 $0.00440$ $0.01555$ $0.05687$ $0.16513$ $0.28418$ $0.77599$ $94.7\%$ $19.3\%$ $0.49029$ $0.03619$ 10000 1.0 $0.00112$ $0.00391$ $0.05606$ $0.08423$ $0.14648$ $0.78001$ $97.7\%$ $98.7\%$ $0.34817$ $0.38824$ 5.0 $0.00011$ $0.00023$ $0.05629$ $0.02701$ $0.03898$ $0.78499$ $98.0\%$ $100.0\%$ $0.16101$ $0.17373$ 0.5 $0.00017$ $0.00219$ $0.05596$ $0.03481$ $0.11630$ $0.78359$ $99.3\%$ $13.7\%$ $0.21625$ $0.01217$ 50000 1.0 $0.00006$ $0.00058$ $0.05587$ $0.01983$ $0.06094$ $0.78346$ $99.7\%$ $100.0\%$ $0.15357$ $0.26129$ 5.0 $0.00002$ $0.00004$ $0.05583$ $0.01062$ $0.01701$ $0.78343$ $99.0\%$ $100.0\%$ $0.07037$ $0.11691$ 0.5 $0.00005$ $0.00134$ $0.05546$ $0.01717$ $0.08593$ $0.78102$ $99.7\%$ $12.3\%$ $0.15604$ $0.00775$ 100000 1.0 $0.00002$ $0.00035$ $0.05555$ $0.01058$ $0.04385$ $0.78172$ $99.7\%$ $0.0\% 0.110$350.11035 $0.21863$ 5.0 $0.00001$ $0.00003$ $0.05563$ $0.00738$ $0.01310$ $0.78230$ $99.7\%$ $100.0\%$ $0.05180$ $0.09784$
@@ 5 Data Analysis
For the PSID data, the point estimators are reasonably accurate, but because the PSID dataset is observational, the results are not as robust as those for the NSW dataset. The non-private CBSR estimate for the PSID data is $0.164$, with a confidence interval from 0.075-0.075 to $0.403$. Under privacy constraints, the estimates, such as the private CBSR with $\epsilon=1.0$, showing $0.035$ with a confidence interval of 0.112-0.112 to $0.182$, reflect greater variability. This variability emphasizes the trade-off between privacy and precision, particularly in observational settings where the underlying data characteristics are less controlled than in experimental settings like the NSW data.
@@
Results for NSW data Results for PSID data Estimator Estimate Confidence Interval Estimate Confidence Interval Non-private CBSR 0.112 $(0.023,0.201)$ 0.164 (0.075,0.403)(-0.075,0.403) Private CBSR ($\epsilon=0.5$) 0.067 (0.004,0.138)(-0.004,0.138) -0.007 (0.147,0.132)(-0.147,0.132) Private CBSR ($\epsilon=1.0$) 0.082 $(0.003,0.162)$ 0.035 (0.112,0.182)(-0.112,0.182) Private CBSR ($\epsilon=5.0$) 0.103 $(0.017,0.188)$ 0.122 (0.030,0.274)(-0.030,0.274)
@@ 6 Concluding Remarks
A promising avenue for future research is the development of an analytical framework for unbounded variables. Our current framework is restricted to bounded variables due to the sensitivity considerations of differential privacy mechanisms. Additionally, the finite-sample performance of our estimators could potentially be improved by more carefully designing the second stage of our privatization process, allowing for a more tailored noise distribution (Awan and Slavković, 2021). Finally, because the proposed methodology targets low- and moderate-dimensional covariates, extending it to high-dimensional settings is important. In particular, the rejection-sampling framework of Awan and Rao (2024) can be computationally burdensome in high dimensions, so more efficient privatization algorithms are needed.
@@ A.1 Proof of Lemma 1
As discussed in Section 2, we consider a scenario where each unit $i$ has an outcome $Y_{i}\in[0,1]$, treatment assignment $Z_{i}\in\{0,1\}$, and covariates $X_{i}\in\mathcal{X}$, where $\mathcal{X}$ is the unit ball so that $\|X_{i}\|_{2}\leq 1$, respectively. Along with the boundedness assumptions, Proposition 2 of Barber and Duchi (2014) implies the minimax risk is lower bounded $\inf_{M_{\epsilon}\in\mathcal{M}_{\epsilon}}\inf_{\hat{\tau}}\sup_{P\in\mathcal{P}}\mathrm{E}\left\{(\hat{\tau}-\tau)^{2}\right\}=\Omega\left\{n^{-1}+(\epsilon n)^{-2}\right\}$. ∎
@@ A.2 Proof of Lemma 2
By Zhao (2019, Theorem 1), for $j=1,\ldots,d$, we have
@@ A.4 Proof of Proposition 3
We first introduce the technical assumptions in Hirano et al. (2003).
@@
The support of $X$ is a Cartesian product of compact intervals. The density of $X$ is bounded, and bounded away from 0.
@@
The convergence of the first component $\xi_{1}$ can be proved by applying the proof of Hirano et al. (2003); Zhao (2019). Under Assumptions 16, Hirano et al. (2003); Zhao (2019) showed that $\hat{\tau}$ is a semiparametric efficient estimator for $\tau$ with the convergence rate $O_{p}(1/\sqrt{n})$. Thus $\xi_{1}$ converges to zero in probability and $\xi_{1}=O_{p}(1/\sqrt{n})$.
[…]
@@
[…]
Under Assumption 1, $\tau^{\prime}(w_{0})$ and $w^{\prime}(\theta_{0})$ are bounded. Additionally, according to Reimherr and Awan (2019, Theorem 3.2), the KNG mechanism ensures that, if the loss function is a twice differentiable convex loss function and the Hessian has strictly positive eigenvalues, we have $\tilde{\theta}-\hat{\theta}=O_{p}(1/np\epsilon).$ Zhao (2019, Proposition 2) showed this condition holds for $\alpha=-1,\beta=0$ or $\alpha=0,\beta=-1$ in the CBSR framework. Therefore, putting these together, we have
@@
$=\{\tau+O_{p}(1/np\epsilon+1/\sqrt{n})+\nu_{1}/a_{n}\}\{1-\nu_{2}/a_{n}+(\nu_{2}/a_{n})^{2}-\cdots\}$